Jekyll Co. has concluded that additional equity financing will be needed to expand operations and that the needed funds will be best obtained through a rights offering. It has correctly determined that as a result of the rights offering, the share price will fall from $67 to $64.00 ($67 is the rights-on price; $64.00 is the ex-rights price, also known as the when-issued price). The company is seeking $19 million in additional funds with a per-share subscription price equal to $34.

How many shares are there currently, before the offering? (Assume that the increment to the market value of the equity equals the gross proceeds from the offering.)
Answer #1

New share = amount raised/subscription price = 19000000/34 = 558823.529

new price = (old price*old shares+amount raised)/(old shares +new shares)

64 = (67*old shares+19000000)/(old shares+558823.529)

64*old shares+558823.529*64 = 67*old shares+19000000

558823.529*64-19000000 = 3*old shares

old shares=5588235.29